OpenStudy (anonymous):

please help 6tanθ−6cotθ=0

OpenStudy (anonymous):

divide both sides by 6 so your equation is now: tanx - cotx = 0 let's use "x" instead of "theta"... easier to type...

OpenStudy (anonymous):

ok....tanx-cotx=0 then wat?

OpenStudy (anonymous):

try writing the equation in terms of only sin and cos....

OpenStudy (anonymous):

ok is 45 135 225 n 315 the solutions?

OpenStudy (anonymous):

idk... i didn't solve.... yet.... maybe these guys will...

OpenStudy (anonymous):

\[\tan \theta = \cot \theta\]\[\frac{ \sin \theta }{ \cos \theta }=\frac{ \cos \theta }{ \sin \theta }\]\[\sin^2 \theta=\cos^2 \theta\]

OpenStudy (anonymous):

\[\sin \theta = \cos \theta\]they only equal the same value at 2 different degrees ( or radians). Are you familiar with the unit circle

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

One answer will be in quadrant 1 the other in 3.

OpenStudy (anonymous):

ok i just got it thank u all!

OpenStudy (anonymous):

hold on. i think you were right above

OpenStudy (anonymous):

if each term is squared then there wil be no negative numbers

OpenStudy (anonymous):

and you will have 4 solutions

OpenStudy (anonymous):

45, 135, 225, 315

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

thank u tho

OpenStudy (tamtoan):

when you have (sin^2)(x) = (cos^2)(x) ....use (cos^2)(x) = 1 - (sin^2)(x) and substitute in, you then can solve because now you only have one variable ... ** i don't know to to write in the form like chme does, or else, it would look much easier.

OpenStudy (anonymous):

its the equation button below the this text field.

OpenStudy (tamtoan):

i got it, thanks @chme very much